The displacement s of a point moving in a straight line is given by: s = 8t 2 + 3t – 5, ‘s’ being in cm and t in s. The initial velocity of the particle is:

  1. The displacement s of a point moving in a straight line is given by:
         s = 8t2 + 3t – 5, ‘s’ being in cm and t in s. The initial velocity of the particle is:

Grade:12

3 Answers

Jheel
6 years ago
By differentiation of the given displacement, We get v=16t+3For initial velocity, u; t=0So, substituting t=0 in 16t+3 We get u=3m/s
Mohd Mujtaba
6 years ago
We know velocity (v)=ds/dt.Here s=8t^2+3t-5.So u= ds/dt =d/dt (8t^2+3t+-5).u=16t+3.At initial velocity t=0.Thus initial velocity =16×0+3=3m/s.Thank.
Harshita kashyap
4 years ago
The x and y coordinate of a particle at any time t are given by: x=7t+4t2 and y=5t where x and y are in m and tin s.the acceleration of the particle at 5sec is _